OpenStudy (anonymous):

how do you simplify 81 to the negative three fourths power?

OpenStudy (bibby):

\(\huge a^{-b}=\frac{1}{a^b}\) \(\huge \sqrt[a]{b^4}={b^4}^\frac{1}{a}=b^{\frac{4}{a}}\)